OpenStudy (anonymous):

A certain arithmetic sequence has the following explicit formula for the nth term: an = 2 + (n - 1)(6) The same sequence has the following recursive formula: an = an-1 + _____ What number belongs in the blank space in the recursive formula?

OpenStudy (john_es):

\[a_{n-1}=2+((n-1)-1)6=2+6(n-1)-6=a_{n}-6\Rightarrow a_n=a_{n-1}+6\]